# 4K4DGen: Panoramic 4D Generation at 4K Resolution

### v2

@@ Abstract
The blooming of virtual reality and augmented reality (VR/AR) technologies has driven an increasing demand for the creation of high-quality, immersive, and dynamic environments. However, existing generative techniques either focus solely on dynamic objects or perform outpainting from a single perspective image, failing to meet the needs of VR/AR applications. In this work, we tackle the challenging task of elevating a single panorama to an immersive 4D experience. For the first time, we demonstrate the capability to generate omnidirectional dynamic scenes with 360-degree views at 4K resolution, thereby providing an immersive user experience. Our method introduces a pipeline that facilitates natural scene animations and optimizes a set of 4D Gaussians using efficient splatting techniques for real-time exploration. To overcome the lack of scene-scale annotated 4D data and models, especially in panoramic formats, we propose a novel Panoramic Denoiser that adapts generic 2D diffusion priors to animate consistently in 360-degree images, transforming them into panoramic videos with dynamic scenes at targeted regions. Subsequently, we elevate the panoramic video into a 4D immersive environment while preserving spatial and temporal consistency. By transferring prior knowledge from 2D models in the perspective domain to the panoramic domain and the 4D lifting with spatial appearance and geometry regularization, we achieve high-quality Panorama-to-4D generation at a resolution of (4096 $\times$ 2048) for the first time. See the project website at https://4k4dgen.github.io.
Figure 1: 4K4DGen accepts a static panoramic image at a resolution of 4096$\times$2048. Upon user interaction through designated clicks, it animates specific regions and converts them into 4D point-based representations. This functionality enables real-time rendering of novel views and dynamic timestamps, thereby enhancing immersive virtual exploration.
@@ 1 Introduction
With the increasing prevalence of generative techniques rombach2022high ; blattmann2023stable , the capability to create high-quality assets has the potential to revolutionize content creation across VR/AR and other spatial computing platforms. Unlike 2D displays such as smartphones or tablets, ideal VR/AR content should offer an immersive and seamless experience, supporting 6-DoF virtual tours, such as high-resolution panoramic 4D environments with dynamic scene elements. Despite significant advancements in the generation of images, videos, and 3D models, the development of panoramic 4D content has lagged, primarily due to the scarcity of well-annotated, high-quality 4D training data. Even in the most relevant field of 4D generation, existing works mainly focus on generating or compositing object-level contents bahmani20234d , which are often in low-resolution (e.g., below 1080p) and cannot fulfill the demand of qualified immersive experiences. Based on these observations, we propose that an ideal generative tool for creating immersive environments should possess the following properties: (i) it should generate content with high perceptual quality, achieving high-resolution (4K) output with dynamic elements (4D); (ii) the 4D representation must be capable of being rendered as coherent, continuous, and seamless 360-degree panoramic views, supporting 6-DoF virtual tours. However, creating diverse 4D panoramic assets of high quality presents two significant challenges: (i) The scarcity of large-scale, annotated 4D data, particularly in panoramic formats, inhibits the training of specialized models. (ii) Achieving both fine-grained local details and global coherence in 4D and 4K panoramic views is challenging for existing 2D diffusion models. These models, typically trained on perspective images with narrow fields of view (FoV), cannot be straightforwardly adapted to the expansive scopes of large panoramic images (see Exp. 4.3). On another front, diffusion models trained with web-scale multi-modal data have demonstrated versatile utility as region-based dynamic priors, and Gaussian Splatting kerbl20233d has shown efficient capabilities in modeling 4D environment. Thus, we formulate the 4D panoramic generation problem by animating a static panoramic image into a video, and subsequently elevate the panoramic video into 4D environment assets. Toward this goal, we focus on and address two bottlenecks: (i) the challenge of animating consistent panoramic videos that maintain object dynamics across a 360-degree field-of-view (FoV) using diffusion models originally trained on narrow FoV perspective images, and (ii) the challenge of robustly preserving spatial and temporal consistency as the panoramic video transitions into a 4D environment.
@@
In this paper, we introduce 4K4DGen, a novel framework designed to facilitate the creation of panoramic 4D environments at resolutions up to 4K. Specifically, we propose the Panoramic Denoiser, which animates specific regions within 360-degree FoV panoramic images by denoising spherical latent codes. This process utilizes a 2D diffusion model initially trained on perspective images with narrow FOV, ensuring global coherence and continuity across the entire panorama. To elevate the omnidirectional panoramic video to a 4D environment, we introduce a 4D lifting mechanism. This mechanism corrects scale discrepancies using a depth estimator enriched with prior knowledge to generate panoramic depth and employs time-dependent 3D Gaussians for cross-frame consistency in dynamic scene representation and rendering. By adapting generic 2D prior models to the panoramic format and effectively lifting 2D dynamics into 4D content with a structured set of 3D Gaussians, we achieve groundbreaking 4K panorama-to-4D content generation. Our contributions can be summarized as follows.
  • We introduce the first framework capable of creating high-resolution (up to 4096$\times$2048) 4D omnidirectional assets. This framework addresses the process through two key techniques: panoramic video animation and 4D panoramic lifting. These techniques enable coherent and seamless 4D panoramic rendering, even in the absence of annotated 4D data.
@@
  • We propose the Panoramic Denoiser that transfers the generative priors from pre-trained 2D perspective diffusion models to panoramic space, which consistently animate panoramas with scene dynamics.
  • We further elevate dynamic panoramic videos into 4D environments using a structured set of 3D Gaussians. Additionally, we have developed a spatial-temporal geometry alignment mechanism that ensures cross-frame consistency and coherence in the 4D environment.
@@ 2 Related Work
Figure 2: Panoramic Denoiser. We propose a method that involves simultaneous denoising of perspective views and their integration into spherical latents at each denoising step. This approach ensures view-consistent animation across different views.
Recent advancements have significantly expanded the capabilities of generating 2D images using diffusion models, as evidenced in several studies dhariwal2021diffusion ; nichol2021glide ; podell2023sdxl ; ramesh2022hierarchical ; saharia2022photorealistic . Notably, Stable Diffusion rombach2022high optimizes diffusion models (DMs) within the latent spaces of autoencoders, striking an effective balance between computational efficiency and high image quality. Beyond text conditioning, there is increasing emphasis on integrating additional control signals for more precise image generation mou2024t2i ; zhang2023adding . For example, ControlNet zhang2023adding enhances the Stable Diffusion encoder to seamlessly incorporate these signals. Furthermore, the generation of multi-view images is gaining attention, with techniques like MVDiffusion Tang2023mvdiffusion or Geometry Guided Diffusion song2023roomdreamertextdriven3dindoor processing perspective images with a pre-trained diffusion model. Diffusion models are also extensively applied in video generation, as demonstrated by various recent works ge2023preserve ; ho2022imagen ; wang2023modelscope ; wu2023tune ; wu2023lamp ; zhou2022magicvideo . For instance, Imagen Video ho2022imagen utilizes a series of video diffusion models to generate videos from textual descriptions. Similarly, Make-A-Video singer2022make advances a diffusion-based text-to-image model to create videos without requiring paired text-video data. MagicVideo zhou2022magicvideo employs frame-wise adaptors and a causal temporal attention module for text-to-video synthesis. Video Latent Diffusion Model (VLDM) blattmann2023align incorporates temporal layers into a 2D diffusion model to generate temporally coherent videos.
@@
In recent 3D computer vision, a large-scale scene is usually represented as implicit or explicit fields for its appearance mildenhall2020nerf ; kerbl20233d , geometry peng2020convolutional ; wang2023alto ; huang2023nksr , and semantics kerr2023lerf ; zhou2024feature ; qin2023langsplat . We mainly discuss the 3D Gaussian Splatting (3DGS) kerbl20233d based generation here. Several works including DreamGaussian tang2023dreamgaussian , GaussianDreamer yi2023gaussiandreamer , GSGEN chen2023text , and CG3D vilesov2023cg3d employ 3DGS to generate diverse 3D objects and lay the foundations for compositionality, while LucidDreamer chung2023luciddreamer , Text2Immersion ouyang2023text2immersion , GALA3D zhou2024gala3d , RealmDreamer shriram2024realmdreamer , and DreamScene360 zhou2024dreamscene360 aim to generate static large-scale 3D scenes from text. Considering the current advancements in 3D generation, investigations into 4D generation using 3DGS representation have also been conducted. DreamGaussian4D ren2023dreamgaussian4d accomplishes 4D generation based on a reference image. AYG ling2023align equips 3DGS with dynamic capabilities through a deformation network for text-to-4D generation. Besides, Efficient4D pan2024fast and 4DGen yin20234dgen explore video-to-4D generation, and utilize SyncDreamer liu2023syncdreamer to produce multi-view images from input frames as pseudo ground truth for training a dynamic 3DGS.
@@
A panorama is an image that captures a wide, unbroken view of an area, typically encompassing a field of vision much wider than what a standard photo would cover, providing a more immersive representation of the subject. Recently, novel view synthesis using panoramic representation has been widely explored. For instance, PERF wang2023perf trains a panoramic neural radiance field from a single panorama to synthesize 360 novel views. 360Roam huang2022360roam proposed learning an omnidirectional neural radiance field and progressively estimating a 3D probabilistic occupancy map to speed up volume rendering. OmniNeRF gu2022omni introduced an end-to-end framework for training NeRF using only 360 RGB images and their approximate poses. PanoHDR-NeRF gera2022casual learns the full HDR radiance field from a low dynamic range (LDR) omnidirectional video by freely moving a standard camera around. In the realm of 3DGS, 360-GS bai2024360 takes 4 panorama images and 2D room layouts as scene priors to reconstruct the panoramic Gaussian radiance field. 4K4Dxu20244k4d extends 3DGS to high-resolution dynamic scene representation. Most related to our 4K4DGen is DreamScene360 zhou2024dreamscene360 , which is a text-to-3D scene generation with a generated 2D panorama as an intermediate global representation. However, it does not consider dynamics. Our work provides solutions to this problem in the following sections.
@@ 3 Methodology
Figure 3: Overall Pipeline. Initially, panoramic videos are generated using the Panoramic Denoiser during the animating phase. Subsequently, these videos are transformed into a spatially and temporally consistent 4D representation during the 4D lifting phase.
Taking a single panoramic image as input, the aim of 4K4DGen is to generate a panoramic 4D environment that allows for the rendering of novel views at arbitrary angles and timestamps. Our approach initially constructs a panoramic video and then elevates it to 4D Gaussians, thereby facilitating efficient splatting for flexible rendering. However, naïvely animating projected perspective images often leads to unnatural motions and inherently inconsistent animations. To address these challenges, our method proposes the denoising of projected spherical latents, thereby enabling consistent animation of the panoramic video from the source image, as detailed in Sec. 3.3. Furthermore, directly converting multiple perspective images from different timestamps into 4D often results in degenerated geometry and noticeable artifacts (see Sec. 4.3); we address this by leveraging spatial-temporal geometry fusion to lift the panoramic video, as described in Sec. 3.4. The entire pipeline of 4K4DGen is illustrated in Fig. 3.
@@ 3.1 Preliminaries
Latent Diffusion Models (LDMs). LDMs rombach2021highresolution consist of a forward procedure $q$ and a backward procedure $p$. The forward procedure gradually introduces noise into the initial latent code $x_{0}\in{}^{h\times w\times c}$, where $x_{0}=\mathcal{E}(I)$ is the latent code of image $I$ within the latent space of a VAE, denoted by $\mathcal{E}$. Given the latent code at step $t-1$, the $q$ procedure is described as $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\boldsymbol{I})$. Conversely, the backward procedure $p$, aimed at progressively removing noise, is defined as $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. In practical applications, images are generated under the condition $y$, by progressively sampling from $x_{T}$ down to $x_{0}$. Recently, image-to-video (I2V) generation has been realized guo2023animatediff ; dai2023animateanything by extending the latent code with an additional frame dimension and performing decoding at each frame. The denoising procedure is succinctly represented as $x_{t-1}=\Phi(x_{t},I)$, where $x_{t},x_{t-1}\in{}^{l\times h\times w\times c}$ represent the sampled latent codes and $I$ the conditioning image.
@@
Panoramic images or videos, denoted as $I$, are represented using equirectangular projections, forming an $H\times W\times C$ matrix where $H$ and $W$ represent the resolution of the image and $C$ is the number of channels. Although this format preserves the matrix structure, aligning with planar images captured by conventional cameras, it introduces distortions, particularly noticeable at the polar regions of the projected image. To address this, we adopt a spherical representation for panoramas, defining pixel values on a sphere $\mathbb{S}^{2}=\{\boldsymbol{d}=(x,y,z)|x,y,z\in\real\wedge|\boldsymbol{d}|=1\}$. For a formal definition of the projection, we employ a mapping $\mathcal{E}_{I}:[-1,1]^{2}\rightarrow{}^{C}$, normalizing the image coordinates into the range $[0,1]$. Therefore, for a given pixel $\left(x,y\right)\in[-1,1]^{2}$, the corresponding pixel value is derived from $\mathcal{E}_{I}\left(x,y\right)$. We define the spherical representation of panoramas using the field $\mathcal{S}_{I}:\mathbb{S}^{2}\rightarrow{}^{C}$, where $\mathcal{S}_{I}\left(\boldsymbol{d}\right)$ gives the pixel value at a given direction $\boldsymbol{d}=(x,y,z)$. The relationship between the spherical and equirectangular representations is established through the following projection formula:
@@
For perspective images, we establish a virtual camera centered at the origin. The rays for all pixels are defined through ray casting, as described in mildenhall2020nerf , where each ray $\boldsymbol{d}$ is represented by $\boldsymbol{r}(x,y,f,\boldsymbol{u},\boldsymbol{s},R)$. This representation considers the focal length $f$, the z-axis direction $\boldsymbol{u}$, the image plane size $\boldsymbol{s}$, and the camera rotation along the z-axis $R$. Consequently, for a given panorama $I$, the perspective image $P$ can be projected using these camera parameters ($f,\boldsymbol{u},\boldsymbol{s},R$) as:
$\mathcal{E}_{P}\left(x,y\right)=\mathcal{S}_{I}\circ\boldsymbol{r}\left(x,y,f,\boldsymbol{u},\boldsymbol{s},R\right).$ (2)
In this paper, we fix the focal length $f$, image plane size $\boldsymbol{s}$, and the rotation $R$. We denote the process of projecting the panorama $I$ into a perspective image $i$, based on the camera’s $z$-axis direction $\boldsymbol{u}$, as $i=\gamma(I,\boldsymbol{u})$.

### v3

# Panoramic 4D Generation at 4K Resolution

@@ Abstract
The blooming of virtual reality and augmented reality (VR/AR) technologies has driven an increasing demand for the creation of high-quality, immersive, and dynamic environments. However, existing generative techniques either focus solely on dynamic objects or perform outpainting from a single perspective image, failing to meet the requirements of VR/AR applications that need free-viewpoint, 360 virtual views where users can move in all directions. In this work, we tackle the challenging task of elevating a single panorama to an immersive 4D experience. For the first time, we demonstrate the capability to generate omnidirectional dynamic scenes with 360 views at 4K (4096 $\times$ 2048) resolution, thereby providing an immersive user experience. Our method introduces a pipeline that facilitates natural scene animations and optimizes a set of dynamic Gaussians using efficient splatting techniques for real-time exploration. To overcome the lack of scene-scale annotated 4D data and models, especially in panoramic formats, we propose a novel Panoramic Denoiser that adapts generic 2D diffusion priors to animate consistently in 360 images, transforming them into panoramic videos with dynamic scenes at targeted regions. Subsequently, we propose Dynamic Panoramic Lifting to elevate the panoramic video into a 4D immersive environment while preserving spatial and temporal consistency. By transferring prior knowledge from 2D models in the perspective domain to the panoramic domain and the 4D lifting with spatial appearance and geometry regularization, we achieve high-quality Panorama-to-4D generation at a resolution of 4K for the first time.
Figure 1: 4K4DGen takes a static panoramic image with a resolution of 4096$\times$2048 and allows animation through user interaction or an input mask, transforming the static panorama into dynamic Gaussian Splatting. 4K4DGen supports the rendering of novel views at various timestamps, enriching immersive virtual exploration.
@@ 1 Introduction
With the increasing growth of generative techniques (Rombach et al., 2022; Blattmann et al., 2023a), the capability to create high-quality assets has the potential to revolutionize content creation across VR/AR and other spatial computing platforms. Unlike 2D displays such as smartphones or tablets, ideal VR/AR content must deliver an immersive and seamless experience, enabling 6-DoF virtual tours and supporting high-resolution 4D environments with omnidirectional viewing capabilities. Despite significant advancements in the generation of images, videos, and 3D models, the development of panoramic 4D content has lagged, primarily due to the scarcity of well-annotated, high-quality 4D training data. Even in the most relevant field of 4D generation, existing works mainly focus on generating or compositing object-level contents (Bahmani et al., 2023; Lin et al., 2024), which are often in low-resolution (e.g., below 1080p) and cannot fulfill the demand of qualified immersive experiences. Based on these observations, we propose that an ideal generative tool for creating immersive environments should possess the following properties: (i) the generated content should exhibit high perceptual quality, reaching high-resolution (4K) output with dynamic elements (4D); (ii) the 4D representation must be capable of rendering coherent, continuous, and seamless 360 panoramic views in real time, supporting efficient 6-DoF virtual tours. However, creating diverse, high-quality 4D panoramic assets presents two significant challenges: (i) the scarcity of large-scale, annotated 4D data, particularly in panoramic formats, limits the training of specialized models. (ii) achieving both fine-grained local details and global coherence in 4D and 4K panoramic views is difficult for existing 2D diffusion models. These models, typically trained on perspective images with narrow fields of view (FoV), cannot be easily adapted to the expansive scopes of large panoramic images (see Exp. 4.3). On another front, video diffusion models (An et al., 2023) trained with web-scale multi-modal data have demonstrated versatile utility as region-based dynamic priors, and Gaussian Splatting (Kerbl et al., 2023) has shown efficient capabilities in modeling 4D environment. Thus, we address the large-scale, omnidirectional dynamic scene generation (4D panoramic generation) problem by utilizing the generative power of diffusion models to animate static panoramic images, transforming them into realistic, dynamic scenes that can support immersive, 360 viewing experiences. To achieve this, we propose to elevate the dynamic panoramic video to 4D environment assets using a set of dynamic Gaussians, which can be seamlessly integrated into VR/AR platforms for real-time rendering and interaction.
@@
In this paper, we introduce 4K4DGen, a novel framework designed to enable the creation of panoramic 4D environments at resolutions up to 4K. 4K4DGen addresses the key challenges of maintaining consistent object dynamics across the entire 360 field-of-view (FoV) in panoramic videos, while preserving both spatial and temporal coherence as the video transitions into a fully interactive 4D environment. Specifically, we propose the Panoramic Denoiser, which animates 360 FoV panoramic images by denoising spherical latent codes corresponding to user-interacted regions. The Panoramic Denoiser leverages a well-trained diffusion model originally designed for narrow-FoV perspective images, enabling the generation of 360 dynamic panoramas while ensuring global coherence and continuity throughout the entire panorama. To transform the omnidirectional panoramic video into a 4D environment, we introduce Dynamic Panoramic Lifting, which corrects scale discrepancies using a depth estimator enriched with perspective prior knowledge to generate panoramic depth maps. Additionally, it employs time-dependent 3D Gaussians optimized with spatial-temporal geometry alignment to ensure cross-frame consistency in dynamic scene representation and rendering. By adapting generic 2D statistical patterns from the perspective domain to the panoramic format and effectively regularizing Gaussian optimization with geometric principles, we achieve high-quality 4K panorama-to-4D content generation with photorealistic novel-view synthesis capabilities. Our contributions can be summarized as follows.
  • We introduce 4K4DGen, the first framework capable of generating high-resolution (up to 4096$\times$2048) 4D omnidirectional assets without the need for annotated 4D data.
@@
  • We propose the Panoramic Denoiser, which transfers generative priors from pre-trained 2D perspective diffusion models to the panoramic space, enabling consistent animation of panoramas with dynamic scene elements.
  • We introduce Dynamic Panoramic Lifting, a method that transforms dynamic panoramic videos into dynamic Gaussians, incorporating spatial-temporal regularization to ensure cross-frame consistency and coherence.
Figure 2: Panoramic Denoiser adapts diffusion priors from the perspective domain to the panoramic domain by simultaneously denoising perspective views and integrating them into spherical latents at each denoising step. This approach ensures consistent animation across multiple views.
@@ 2 Related Work
Recent advancements have significantly expanded the capabilities of generating 2D images using diffusion models, as evidenced in several studies (Dhariwal & Nichol, 2021; Nichol et al., 2021; Podell et al., 2023; Ramesh et al., 2022; Saharia et al., 2022). Notably, Stable Diffusion (Rombach et al., 2022) optimizes diffusion models (DMs) within the latent spaces of autoencoders, striking an effective balance between computational efficiency and high image quality. Beyond text conditioning, there is increasing emphasis on integrating additional control signals for more precise image generation (Mou et al., 2023; Zhang et al., 2023). For example, ControlNet (Zhang et al., 2023) enhances the Stable Diffusion encoder to seamlessly incorporate these signals. Furthermore, the generation of multi-view images is gaining attention, with techniques like MVDiffusion (Tang et al., 2023b) or Geometry Guided Diffusion (Song et al., 2023) processing perspective images with a pre-trained diffusion model. Diffusion models are also extensively applied in video generation, as demonstrated by various recent works (Ge et al., 2023; Ho et al., 2022; Wang et al., 2023b; Wu et al., 2023c; d; Zhou et al., 2022). For instance, Imagen Video (Ho et al., 2022) utilizes a series of video diffusion models to generate videos from textual descriptions. Similarly, Make-A-Video (Singer et al., 2022) advances a diffusion-based text-to-image model to create videos without requiring paired text-video data. MagicVideo (Zhou et al., 2022) employs frame-wise adaptors and a causal temporal attention module for text-to-video synthesis. Video Latent Diffusion Model (VLDM) (Blattmann et al., 2023b) incorporates temporal layers into a 2D diffusion model to generate temporally coherent videos.
@@
In recent 3D computer vision, a large-scale scene is usually represented as implicit or explicit fields for its appearance (Mildenhall et al., 2020; Kerbl et al., 2023), geometry (Peng et al., 2020; Wang et al., 2023c; Huang et al., 2023), and semantics (Kerr et al., 2023; Zhou et al., 2024a; Qin et al., 2023). We mainly discuss the 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) based generation here. Several works including DreamGaussian (Tang et al., 2023a), GaussianDreamer (Yi et al., 2023), GSGEN (Chen et al., 2023), and CG3D (Vilesov et al., 2023) employ 3DGS to generate diverse 3D objects and lay the foundations for compositionality, while LucidDreamer (Chung et al., 2023), Text2Immersion (Ouyang et al., 2023), GALA3D (Zhou et al., 2024c), RealmDreamer (Shriram et al., 2024), and DreamScene360 (Zhou et al., 2024b) aim to generate static large-scale 3D scenes from text. Considering the current advancements in 3D generation, investigations into 4D generation using 3DGS representation have also been conducted. DreamGaussian4D (Ren et al., 2023) accomplishes 4D generation based on a reference image. AYG (Ling et al., 2023) equips 3DGS with dynamic capabilities through a deformation network for text-to-4D generation. Besides, Efficient4D (Pan et al., 2024) and 4DGen (Yin et al., 2023) explore video-to-4D generation, and utilize SyncDreamer (Liu et al., 2023) to produce multi-view images from input frames as pseudo ground truth for training a dynamic 3DGS. 4K4D (Xu et al., 2024) is a high-resolution reconstruction technique that extends 3DGS to model complex human motion with detailed backgrounds while achieving real-time rendering speed.
@@
A panorama is an image that captures a wide, unbroken view of an area, typically encompassing a field of vision much wider than what a standard photo would cover, providing a more immersive representation of the subject. Recently, novel view synthesis using panoramic representation has been widely explored. For instance, PERF (Wang et al., 2023a) trains a panoramic neural radiance field from a single panorama to synthesize 360 novel views. 360Roam (Huang et al., 2022) proposed learning an omnidirectional neural radiance field and progressively estimating a 3D probabilistic occupancy map to speed up volume rendering. OmniNeRF (Gu et al., 2022) introduced an end-to-end framework for training NeRF using only 360 RGB images and their approximate poses. PanoHDR-NeRF (Gera et al., 2022) learns the full HDR radiance field from a low dynamic range (LDR) omnidirectional video by freely moving a standard camera around. In the realm of 3DGS, 360-GS (Bai et al., 2024) takes 4 panorama images and 2D room layouts as scene priors to reconstruct the panoramic Gaussian radiance field. DreamScene360 (Zhou et al., 2024b) achieves text-to-3D Panoramic Gaussian Splatting by utilizing monocular depth priors to regularize the Gaussian optimization.
@@ 3 Methodology
Taking a single panoramic image as input, the goal of 4K4DGen is to generate a panoramic 4D environment capable of rendering novel views from arbitrary angles and at various timestamps. Our approach initially constructs a panoramic video and then elevates it into a series of 3D Gaussians, enabling efficient splatting for flexible rendering. Naïvely animating projected perspective images, however, often results in unnatural motion and inconsistent animations. To overcome this, our method propose the denoising of projected spherical latents, ensuring consistent animation of the panoramic video from the original image, as detailed in Sec. 3.3.
Moreover, directly converting multiple perspective images from different timestamps into 4D frequently leads to degraded geometry and visible artifacts (see Sec. 4.3). We address this by applying spatial-temporal geometry fusion to lift the panoramic video, as described in Sec. 3.4. The complete pipeline of 4K4DGen is illustrated in Fig. 3.
Figure 3: Overall Pipeline. Beginning with a static panorama as input, the Animating Phase generates a panoramic video by first mapping the panorama into a spherical latent space, followed by denoising within the perspective space, fusing back to the spherical latent space at each step, and finally transforming it into the panoramic space. In the 4D Lifting Phase, a series of dynamic Gaussians is employed to lift the panoramic video into a 4D representation, ensuring both spatial and temporal consistency.
@@ 3.1 Preliminaries
Latent Diffusion Models (LDMs).
LDMs  (Rombach et al., 2021) consist of a forward procedure $q$ and a backward procedure $p$. The forward procedure gradually introduces noise into the initial latent code $x_{0}\in{}^{h\times w\times c}$, where $x_{0}=\mathcal{E}(I)$ is the latent code of image $I$ within the latent space of a VAE, denoted by $\mathcal{E}$. Given the latent code at step $t-1$, the $q$ procedure is described as $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\bm{I})$. Conversely, the backward procedure $p$, aimed at progressively removing noise, is defined as $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. In practical applications, images are generated under the condition $y$, by progressively sampling from $x_{T}$ down to $x_{0}$. Recently, image-to-video (I2V) generation has been realized  (Guo et al., 2023; Dai et al., 2023) by extending the latent code with an additional frame dimension and performing decoding at each frame. The denoising procedure is succinctly represented as $x_{t-1}=\Phi(x_{t},I)$, where $x_{t},x_{t-1}\in{}^{l\times h\times w\times c}$ represent the sampled latent codes and $I$ the conditioning image. Recently, image-to-video (I2V) generation has been achieved (Guo et al., 2023; Dai et al., 2023) by extending the latent code with an additional frame dimension and performing decoding at each frame. The denoising procedure is succinctly expressed as $x_{t-1}=\Phi(x_{t},I)$, where $x_{t},x_{t-1}\in{}^{l\times h\times w\times c}$ represent the sampled latent codes, and $I$ represents the conditioning image.
@@
Panoramic images or videos, denoted as $I$, are typically represented using equirectangular projections, forming an $H\times W\times C$ matrix, where $H$ and $W$ denote the image resolution and $C$ represents the number of channels. While this format preserves the matrix structure, making it consistent with planar images captured by conventional cameras, it introduces distortions, especially noticeable near the polar regions of the projection. To mitigate these distortions, we adopt a spherical representation for panoramas, where pixel values are defined on a sphere $\mathbb{S}^{2}=\{\bm{d}=(x,y,z)|x,y,z\in\real\wedge|\bm{d}|=1\}$. For a more precise definition of the projection, we represent matrix-like images using a mapping $\mathcal{E}_{I}:[-1,1]^{2}\rightarrow{}^{C}$, which normalizes the image coordinates into the range $[0,1]$. Thus, for any given pixel $\left(x,y\right)\in[-1,1]^{2}$, the corresponding pixel value is determined by $\mathcal{E}_{I}\left(x,y\right)$. We define the spherical representation of panoramas using the field $\mathcal{S}_{I}:\mathbb{S}^{2}\rightarrow{}^{C}$, where $\mathcal{S}_{I}\left(\bm{d}\right)$ gives the pixel value at a given direction $\bm{d}=(x,y,z)$. The relationship between the spherical and equirectangular representations is established through the following projection formula:
@@
For perspective images, we define a virtual camera centered at the origin. The rays for each pixel are determined through ray casting, as described in  (Mildenhall et al., 2020), where each ray $\bm{d}$ is represented by $\bm{r}(x,y,f,\bm{u},\bm{s},R)$. This representation takes into account the focal length $f$, the z-axis direction $\bm{u}$, the image plane size $\bm{s}$, and the camera’s rotation along the z-axis $R$. Consequently, for a given panorama $I$, the perspective image $P$ can be projected using these camera parameters ($f,\bm{u},\bm{s},R$) as:
$\mathcal{E}_{P}\left(x,y\right)=\mathcal{S}_{I}\circ\bm{r}\left(x,y,f,\bm{u},\bm{s},R\right).$ (2)
In this paper, we fix the focal length $f$, the image plane size $\bm{s}$, and the rotation $R$. We denote the process of projecting the panorama $I$ into a perspective image $i$, based on the camera’s $z$-axis direction $\bm{u}$, as $i=\gamma(I,\bm{u})$.